OPTICAL PUMPING AND ELECTRON SPIN RESONANCE OF SINGLE ^{87}Rb ATOMS ON HELIUM NAN-ODROPLETS

\underline{MARKUS KOCH}, JOHANNES POMS, ALEXANDER VOLK, and WOLFGANG E. ERNST, Institute of Experimental Physics, TU Graz, Petersgasse 16, 8010 Graz, Austria.

Our recent development of electron spin resonance (ESR) spectroscopy on superfluid helium nanodroplets (He_{N}) provides a sensitive tool to investigate interactions between a surface located alkali-metal atom and an ESR silent species inside the droplet^{a}. Highest sensitivity is expected for alkali-metal atoms with large hyperfine coupling. We present hyperfine resolved ESR spectra of single ^{87}Rb (hyperfine constant a_{HFS} = 3417 MHz) atoms isolated on He_{N}^{b}. In accordance with our previous work on ^{85}Rb (a_{HFS} = 1012 MHz) we find a droplet size dependent increase of a_{HFS} between 400 and 450 ppm, due to the electronic perturbation by the helium environment. The process of optical pumping and of optical detection on He_{N} is investigated in detail in order to optimize the ESR signal. A simple model for optical pumping on He_{N} is presented, which agrees well with the experimental results.

aM. Koch, G. Auböck, C. Callegari, and W.E. Ernst, Phys. Rev. Lett. 103, 035302 (2009)

bA. Volk, J. Poms, M. Koch, and W.E. Ernst, J. Phys. Chem. A, in press